OpenStudy (anonymous):

The rate constant of a reaction(K) is 0.0047 at 298k and the activation energy is 33.6 Kj/mol. What is K at 348K? T1- 298 T2- 348 K1- 0.0047 K2-? Ea- 33.6 Kj/mol I keep getting 1.01 S^-1 but the textbook says its 0.033 s^-1

OpenStudy (anonymous):

Ln*K2 =0.0092 is as far as I can get.

OpenStudy (aaronq):

are you using the two-point form Arrhenius equation?

OpenStudy (anonymous):

Ln (k2/k1) = -Ea/r((1/t2)-(1/t1)) This is the equation that I'm using

OpenStudy (aaronq):

\(\large ln\dfrac{k_2}{k_1}=-\dfrac{E_a}{R}(\dfrac{1}{T_2}-\dfrac{1}{T_1})\) raise both sides to e \(\large \dfrac{k_2}{k_1}=e^{(\small-\dfrac{E_a}{R}(\dfrac{1}{T_2}-\dfrac{1}{T_1}))}\) move \(k_1\) over \(\large k_2=k_1e^{(\small-\dfrac{E_a}{R}(\dfrac{1}{T_2}-\dfrac{1}{T_1}))}\)

OpenStudy (anonymous):

Ah! I get it now! I was multiplying K1 to the other side and then taking the anti-log of that and that's how I got 1.01. Thanks for the help!

OpenStudy (aaronq):

yeah, logarithms are funny to work with, good stuff though! no problem